OpenStudy (anonymous):

Mr.Bob is playing a game.The rule of the game is the following .Bob has first to throw a die.When the die appears with number X, Bob has chance to open X boxes among 10. Among the 10 boxes, there are two boxes containing the prize of $500, one box containing $1000, and the other boxes are empty. If he gets the prize boxes, he gets the prize. a.Find the probability for him to win the prize. b.How much would he expect to win the prize. c.Suppose that the die thrown by Bob appeared with number 4, and that he has opened three empty boxes already.How much would he expect to win the prize?

OpenStudy (anonymous):

can you tell me how can i select 4 boxes out of 10 so that i can get a prize

OpenStudy (anonymous):

hello....is there anyone???

OpenStudy (anonymous):

yeah is i get to choose 4 out of 10 boxes the probability i get at least one prize is \(1-\frac{7}{10}\times \frac{6}{9}\times \frac{5}{8}\times \frac{4}{7}\)

OpenStudy (anonymous):

and of course the probability i get to choose 4 boxes is \(\frac{1}{6}\)

OpenStudy (anonymous):

which gives a good idea of what a general solution might look like.

OpenStudy (anonymous):

sry but you might be wrong

OpenStudy (anonymous):

You are so clever,but could you give the probability for him to win the prize. Thank in advance for your help.